A Bibliography of Publications in IEEE Computer: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

21 December 2023 Version 2.19

Title word cross-reference

3 [103].

-D [103].

1 [174, 324]. 11 [45]. 1980s [196]. 1987 [193]. 1990s [83].

2D [330].

3 [104, 331]. 3-D [104, 331]. 38 [142]. 3D [183, 330].

5 [277].

754 [26]. 79 [32].

Abingdon [323]. Abstractions [125]. Access [231, 247]. accidents [165]. According [240]. Accuracy [293]. ACM [221]. Acquisition [330]. across [286]. Ada [31, 139, 147, 287]. Adaptable [17, 90]. Adaptive [179, 232, 236, 318]. Added [113]. Addressable [326]. ADMS [153]. Advanced [99, 280, 310]. aerodynamics [337]. Aided [206]. Alamos [334]. Algebra [175]. Algorithm [361]. Algorithmic [133, 291, 316]. Algorithms [8, 49, 105, 109, 157, 178, 265]. Allocation [19]. Alternatives [326]. Ambler [314]. Amdahl [264]. Ames [337]. Analog [234]. Analysis [25, 41, 77, 123]. Animating [122]. Announcements [204, 205]. Appendum [267]. Application [249]. Applications [26, 67, 73, 110, 256, 336]. Applying [64]. Appreciation [239, 267]. Apprentice [268].

Approach [66, 72, 79, 140, 346, 349]. Approaches [3, 250]. April [267]. Arbitrary [265]. Architecture [14, 34, 73, 88, 116, 129, 162, 163, 168, 230, 256, 259, 264, 282]. Architecture/Compiler [168]. Architectures [17, 50, 71, 134, 158, 160, 271]. area [12]. Arithmetic [2, 23, 24, 26, 51]. Armstrong [249]. Array [34–36, 173, 176, 178, 179, 265]. Arrays [135, 157, 172, 177, 181, 182, 243]. ART [236]. Artificial [158, 160]. assembly [296]. Assistant [143, 304]. Associative [326]. Assuring [342]. Astronautics [324]. Augmented [170]. Authority [231]. automated [296]. Automatic [62, 257, 302]. Automation [30, 290, 295–298]. automation-robotics [298]. Autonomous [313–315]. Avalon [275]. Avalon/C [275].

Back [3, 4]. Back-End [3, 4]. Balancing [58]. band [12]. Base [39, 40, 64, 162, 215]. Based [6, 58, 81, 95, 128, 141, 156, 183–188, 301, 359]. batch [296]. Beamforming [179]. Behavior [80]. benchmark [323]. Beyond [124]. Bibliography [238]. Bidirectional [241]. Binary [23, 234, 240]. Birth [279]. Book [127, 211]. Box [248]. Box-Structured [248]. Branch [85]. Broadband [350]. Broadcast [8]. Browsing [140, 225]. Buffer [45, 85]. Building [81, 166, 321]. bullet [165]. Bus [16]. Business [30].

C [275]. C2 [278]. Cabinet [247]. caches [273]. CAD [105, 161, 183, 184, 187, 307]. CAD-based [183]. calculating [9]. Calendar [208]. Call [209]. CAM [105]. Camera [112]. Capacity [119, 142]. Career [210]. Case [197, 214, 273]. CCD [242]. Cellular [88, 180]. Center [336, 337]. Chaining [341]. Characterization [254]. Charlotte [343]. Checking [7]. Chief [192]. Chinese [106]. Chinese/Kanji [106]. Chip [32, 241]. Chips [1]. Circuit [349]. Circuits [7, 241, 242, 302]. Cm* [68]. Code [286]. Codes [7]. coding [362]. Coherence [227]. Collapsed [303]. Collapsed-Fault [303]. Color [78]. Combinatorial [117]. Combining [308]. Commercial [166, 231]. Communication [12, 66]. Comparison [240]. Compiler [11, 168, 287]. Compiler-Compiler [11]. Complete [152, 365]. Complex [197]. Complexity [124]. Components [137]. Composing [284]. Compression [91, 361, 362, 365]. Computation [15, 49, 245, 293]. computational [337]. Computations [346]. Computer [22, 38, 55, 58, 63, 66, 71, 74, 92, 98, 100, 102, 112, 116, 128, 158, 162, 168, 174, 191, 193, 201, 206, 213, 221, 271, 280, 289, 353, 357, 359]. Computer-Aided [206]. Computer-Based [128]. Computers [124, 196, 255, 260, 327]. Computing [12, 35, 48, 51, 149, 221, 234, 237, 263, 288, 353]. Concept [172, 173, 182, 226]. Concepts [110, 261, 310]. Concurrent [51, 108, 255, 346]. Conference [193, 206, 262]. Conferences [207]. Conferencing [128]. Configuration [113, 348]. Configurations [138]. CONLAN [110]. Connection [256]. Connectionist [160]. Considerations [6, 61, 156, 281]. Constrained [319]. Constraints [360]. Construction [104, 226]. Content [326]. Content-Addressable [326]. Contention [269]. contents [190]. Control [89, 144, 185, 188, 231]. Controller [179]. Controlling [318]. Controversy [124]. Convex [278]. Core [221]. Correcting [7]. Correctness [252, 270]. Coupled [229]. cover [189]. Coverage [303]. Covered [152]. Cray [279, 341]. Criteria [240]. Cross [323]. Crossbar [171]. Current [120, 129, 241, 252, 295–298]. Current-Mode [241]. Curriculum [344]. Curved [69]. Cut [113]. Cydra [277].

D [103, 104, 331]. Dallas [193]. DARPA [74, 131]. Data [18, 19, 39–42, 53–56, 59, 64, 91, 105, 106, 162, 182, 215, 250, 265, 329– 332, 338, 356, 359, 361, 365]. data-base [40]. Data-Driven [182, 265]. Data/Knowledge [162, 215]. Database [71, 86, 90, 97, 102, 136, 140, 355, 357]. Databases [126, 151, 155, 161, 214, 360]. Dataflow [52, 133]. Deadlock [13, 351]. Debugging [347]. decisions [277]. Declarative [346]. Decomposing [105]. defense [283]. Definitive [364]. Denormalized [27]. departmental [277]. Depth [321]. Description [111]. Design [1, 6, 45, 66, 85, 89, 90, 108, 125, 136, 139, 144, 145, 153, 156, 179, 198, 206, 216, 253, 272, 277, 278, 281, 304, 325, 349, 359]. Designing [343]. Designs [301]. detection [351]. Determining [78]. Development [10, 121, 217, 222, 246, 301, 305]. Developments [176]. Diagnostic [97]. Digital [80, 230]. direct [273]. direct-mapped [273]. Directed [97]. Directions [14]. Discipline [221, 288, 289]. Disk [59, 119, 156]. Disk-Based [156]. Disks [247]. Display [329]. Distributed [8, 12, 19, 67, 132, 139, 169, 274, 320, 334, 351]. distributions [9]. do [164]. Document [129, 222, 223]. domain [264]. Domesticating [148]. Down [104]. drawing [5]. Drawings [104]. Driven [182, 253, 265]. Dual [184]. Dynamic [266, 324].

Earth [332]. easy [5]. easy-to-use [5]. Editing [223]. Editor [169, 192, 238]. Editor-in-Chief [192]. Editors [84, 218, 355]. education [289]. Efficient [108, 152]. Electrical [213]. Electronic [76, 127, 297, 333]. Electronics [213]. End [3, 4]. Engineer [304]. Engineering [104, 126, 165, 278, 308, 328, 344]. Engineers [213]. Enhancement [246]. Enhancing [126]. Environment [28, 29, 31, 94, 101, 140, 146, 220, 226, 245, 304, 321]. Environments [216, 217, 219, 345]. Errata [24]. Error [7]. Error-Correcting [7]. essence [165]. Essential [114, 178]. Ethernet [63]. Evaluating [138]. Evaluation [166]. Event [227]. Evolution [63, 306, 345, 353]. Exception [294]. Exchange [170]. Execution [46]. Exercise [253]. Expanding [328]. Experience [81, 343]. Experimental [131, 338]. Experiments [62]. Expert [97, 142, 144–146, 166, 261]. Exploration [314]. Expressing [109].

Facilitate [36]. Facility [245, 310, 343].

Facts [225]. Fall [193]. Fault

[96, 177, 229, 303]. Fault-Tolerant [96, 229]. Federal [115]. Field [329]. Fifth [116]. Fifth-Generation [116]. File [62, 162, 247]. Filtering [232]. Finding [225]. Fixed [157]. Fixed-Size [157]. Flames [330]. Flexible [67]. Flight [310]. Floating [2, 23–26]. Floating-Point [2, 23, 25]. Flow [18, 53–56, 64, 329]. Flows [330]. Fluid [329]. FolioPub [224]. Force [221]. Formal [198, 252]. Formalization [219]. Format [363]. Formats [129]. Fortran [36, 251]. Framework [307, 316]. Free [60]. Friends [148]. Front [189]. Functional [79, 150]. Functionally [34]. Fuzzy [261].

GaAs [168, 179]. Gateway [47]. Gateways [73, 74]. General [174, 281]. General-Purpose [174, 281]. Generating [303]. Generation [20, 116, 263, 300, 302]. Genesis [102]. Gentest [302]. geographic [41, 356]. GEOMATIC [103]. Geometric [293]. Good [342]. Government [115]. Graduation [312]. graph [5]. Graphic [103, 363]. Graphical [101, 120, 121, 274, 348]. Graphics [22, 98, 130, 259]. graphs [54]. Grasping [295]. Grids [317]. Guest [84, 169, 238, 355]. Guide [2, 24, 364].

Handling [145, 294]. Hardware [80, 107, 109, 111, 252, 304, 354]. Hashing [266]. Heterogeneous [74, 86, 245]. High [91, 119, 230, 260, 300]. High-Level [300]. High-Performance [91]. High-Speed [119, 230, 260]. Highly [48, 51]. Holy [37]. Hyper [250]. Hyper-Semantic [250]. Hyperchannel [75]. Hypertext [194, 225].

IBM [142, 251, 339, 341]. IC [204]. ICs [240]. IEEE [26, 206]. IEEE-754 [26]. If [60]. Image [71, 355, 358]. impact [12]. Implementation [2, 24, 152, 172, 173, 182, 233]. implementations [110]. Improving [195]. Inc. [213]. individual [362]. Influence [345]. Infomart [193]. Information [41, 42, 130, 226, 272, 321, 356]. Inheritance [275]. Inspection [322]. inspections [270]. Institute [213, 312, 344]. Instruction [93, 124, 197, 324]. Instruction-Level [93]. Integrated [71, 216, 359]. Integrating [35, 136, 181]. Integration [89]. Integrity [360]. Intelligence [158, 160]. Intelligent [47, 295, 313]. Interactive [98, 101, 331]. Interchange [129, 130]. Interconnecting [74, 76, 86, 137]. Interconnection [43, 44, 169, 170, 285]. Interface [125, 309]. Interfaces [274, 284]. Interfacing [105]. Interlisp [29]. Intermedia [226]. Internal [266]. International [129, 206]. Internet [74]. internetworking [73]. Interoffice [132]. Interpreter [70]. Interpreting [33]. InterViews [284]. Introduction [84, 169, 194, 238, 252, 254, 355]. investigate [270]. IS-A [77]. ISDN [350]. Isn’t [77]. ISR [357]. Issues [4, 14, 114, 139, 287].

Joint [193]. Jukebox [119].

Kanji [106]. Kernel [320]. Kingdom [176]. Knowledge [79, 81, 82, 97, 126, 144, 162, 215, 225, 308, 359]. Knowledge-Based [81, 359].

Knowledge-Directed [97]. KRYPTON [79].

Laboratory [334, 335, 338]. LANDSAT [41]. Language [10, 14, 30, 108, 109, 111, 218, 345, 359]. Languages [40, 53, 56]. Large [15, 49, 119, 135, 162, 214, 332]. Large-Capacity [119]. Large-Scale [15]. Lawrence [335]. learned [316]. Learning [315, 319]. length [9]. Level [93, 292, 300]. Library [301]. Library-Based [301]. Linda [148]. Linear [175]. Links [75, 77]. LISP [32, 163, 254]. Literature [212]. Livermore [335]. Load [58]. Local [12, 58, 63]. Localization [187]. Logic [238, 239, 241, 243, 267].

Machine [116, 256]. Machines [56, 64, 163, 215, 313]. Mail [76, 131, 132]. Maintenance [325]. Management [57, 71, 86, 154, 224, 262, 309, 355, 359]. Managing [161]. Manipulating [42]. Manipulation [322]. Manipulators [188]. Manufacturing [187]. Map [356]. mapped [273]. Mapping [157, 178, 272]. Mass [60, 61, 118]. Master [344]. Matrix [51, 174]. Matrix-1 [174]. matures [289]. Maximum [251]. Measuring [9, 69]. Mechanical [104, 144, 296]. Medical [331]. Member [213]. Memories [96]. Memory [38, 57, 269, 326]. Message [191, 192, 255]. Message-Passing [255]. Metadata [154]. Methodology [89, 90]. Methods [198, 327]. Microcomputer [16]. Microprocessor [6, 87, 281]. Microprocessor-Based [6]. Microsystem [205]. Mid [196]. Mid-1980s [196]. Migration [62, 343]. Mini [164]. Mini-Super [164]. Misconceptions [263]. MML [87]. Mobile [315, 317]. Mode [241]. Model [118, 186–188, 233, 246, 321, 358]. Model-Based [186–188]. Modeling [93, 105, 161, 244, 250]. Modular [43]. Monitoring [46]. Motion [33, 234, 291].

Motions [292]. MP [279, 341]. Multicomputer [138, 244]. Multicomputers [255]. Multilevel [73]. Multimedia [131, 156]. Multiple [87, 222, 238, 239, 241–243, 264, 267]. multiple-domain [264]. Multiple-Microprocessor [87]. Multiple-Valued [238, 239, 241–243, 267]. Multiplier [241]. Multiprocessing [117]. Multiprocessor [95, 114, 229, 249, 285]. Multiprocessors [133, 227, 269]. Multisensor [320]. Multistage [170]. Multivalued [240]. Myths [257].

NASA [337]. National [262, 334–336]. Nature [166]. Naval [338]. Navigation [315–317]. Nets [232]. Network [3, 58, 63, 66, 75, 228, 233, 235, 236, 311, 350, 363]. Networks [4, 21, 43, 44, 49, 74, 77, 92, 169– 171, 234, 237, 285]. Neural [232, 233, 235–237, 319]. News [201]. Next [263]. Next-Generation [263]. No [165]. Nonlinear [319]. Null [179]. Numbers [27].

OBE [30]. Object [104, 156, 183, 307, 342, 358]. Object-Oriented [307, 342]. Objects [33, 155]. Occluded [186]. Occupancy [317]. October [193]. Office [30, 129]. Offs [45, 277]. Once [247]. Operating [57, 99, 139, 228]. Operator [143]. Opinion [56]. Opportunities [210, 328]. Optical [59, 119, 156, 171, 215, 247, 350]. Optimization [243, 254, 286]. Optimum [230]. Ordering [227]. Organization [45]. Organizing [236]. Oriented [307, 342]. Overlapping [112]. Overview [11, 13, 198, 268].

package [5]. Packet [43]. Paper [145]. Papers [209]. Para [150]. Para-Functional [150]. Paradigm [83]. Parallel [34, 48, 49, 94, 115, 116, 134, 147, 149, 169, 175, 214, 230, 327]. Parallelism [148, 167].

PARET [244]. Part [78]. Partially [186]. Partitioning [157, 178]. Parts [112, 186, 187]. Passing [255]. Patrec [97]. Pattern [100, 232, 235, 236]. PC [339]. peace [37]. Perception [317]. Performance [4, 91, 95, 123, 133, 138, 156, 251, 285, 347]. Persistent [72]. Perspective [31, 353]. Perspectives [15]. philosophies [277]. Pick [292]. Pick-and-Place [292]. PICT [101]. Pictorial [39–42, 360]. Picture [40]. Pipelining [341]. pivotal [283]. Place [292]. Planetary [314]. Planning [142, 186, 291, 292, 296]. Plausibility [253]. Plausibility-Driven [253]. plea [37]. PNG [363, 364]. Point [2, 23–26]. Poker [94]. Portable [363]. Power [143]. Practical [55, 220]. Precision [282]. Precompilation [36]. Predictable [320]. Prediction [85, 185]. Prediction-Based [185]. Preliminary [221]. President [191]. Price [113, 159]. Pride [145]. Principles [153]. Problem [13, 263]. Problems [117, 293]. procedures [286]. Process [89, 343]. Processing [19, 36, 41, 51, 71, 106, 147, 158, 169, 180, 218, 229, 230, 299, 356, 357]. Processor [35, 93, 113, 115, 135, 175, 182, 254]. Processors [34, 173, 178, 196]. Product [202]. Production [11, 141]. Productivity [195]. products [203]. Program [46, 54, 93, 121, 198, 218, 270]. Programmable [243, 311]. Programmer [268]. Programming [28, 29, 72, 87, 94, 101, 120, 150, 167, 188, 216, 219, 220, 251, 257, 296, 348, 359]. Programs [36, 115, 122, 342]. project [11, 110, 336]. PROLOG [108]. Properties [275]. Proposals [25]. Proposed [2, 16, 23, 24, 26]. Prospects [257]. Protocols [350]. Prototypes [309]. Prototyping [297, 305, 306, 308, 310, 311]. PSN [81]. Publication [224]. Purpose [1, 174, 281].

QBE [30]. QBE/OBE [30]. Quality [11]. Quantum [333]. Query [40, 299]. Query-Processing [299]. Querying [214]. queue [9].

R&D [262]. RAM [326]. Random [247]. Random-Access [247]. Rapid [305, 306, 308, 310]. rate [362]. Reaching [22]. Reading [127]. Real [67, 128, 230, 263, 320]. Real-Time [67, 128, 230, 263, 320]. Recognition [100, 112, 232, 235, 236, 358]. Recognizing [186]. Recommendations [344]. Reconfiguration [134]. Recovery [275, 325]. Reduced [254]. Reduced-Introduction-Set [254]. Reducing [269]. Reduction [95]. Reduction-Based [95]. Reference [118, 365]. Refinement [248]. Relationships [78]. Relief [103]. Remote [245]. report [221, 289, 352, 354]. Representation [79, 126, 183, 222, 272, 329, 330]. Representations [184]. Representing [80, 82]. Research [116, 252, 268, 295–298, 310, 328, 337, 338]. Resistive [234]. retrospective [340]. Reuse [325]. Reusing [137]. Reviews [202, 211]. RISC [65, 196]. Rivers [336]. Robot [69, 183, 185, 188, 292, 315–317]. Robotic [296, 322]. Robotics [290, 291, 294–298]. Robustness [293]. Rochester [47]. Role [130, 333]. Routing [92]. Rover [314]. Rule [141]. Rule-Based [141].

Saxpy [174]. Scale [15, 49]. Schedule [141]. Scheduling [324]. Scheme [32]. Scheme-79 [32]. science [221, 289]. Sciences [332]. Scientific [15, 35, 220, 280, 328, 336, 337]. SDI [146]. Seamless [226]. Search [117]. Second [20, 56]. Secure [299]. Segue [274]. Self [7, 236]. Self-Checking [7].

Self-Organizing [236]. Semantic [77, 250]. Semiconductor [96]. Semiconductors [333]. sequences [362]. Sequential [302, 361]. Sequoia [229]. Serious [263]. Server [156]. Set [87, 254, 303]. Sets [124, 197, 329, 332]. Shapes [104]. Shared [269]. Shared-Memory [269]. Shuffle [170]. Shuffle-Exchange [170]. Shuffles [44]. Signal [51, 180, 230]. silver [165]. Simulated [95]. Simulation [8, 103, 123, 333]. Single [112, 303]. Single-Stuck-Fault [303]. Size [157]. SLAPP [175]. Smalltalk [122, 159]. Snowbird [289]. Society [191, 201, 213]. Software [61, 83, 121, 137, 165, 195, 217, 246, 249, 276, 283, 305, 306, 308, 344]. Solid [105]. Some [176, 198]. Source [231]. Soviet [260]. Space [33]. Spatial [360]. Special [1]. Special-Purpose [1]. Specification [360]. spectrum [353]. Speed [119, 230, 260, 341]. Spherical [184]. Spiral [246]. Sprite [228]. Stack [164]. Standard [2, 16, 23–26, 309]. Standardization [129]. Standards [130, 199]. Star [340]. Static [266]. Status [129]. Steering [179]. Step [178]. Stepwise [248]. Storage [4, 59–61, 118]. Storming [308]. Strategic [262, 283]. Strategies [85, 92, 134, 299]. Strategy [186]. Structure [80]. Structured [237, 248, 349]. Structures [42, 51, 105]. Stuck [303]. Study [214]. Style [342]. Styles [223]. summary [354]. Super [17, 164]. Supercomputer [259, 277, 279, 354]. Supercomputers [18, 20, 276]. Supercomputing [336]. Supersystem [181]. Support [89, 121, 274, 304]. Surfaces [69]. Surrogate [162]. Survey [46, 48, 120, 130, 194, 196, 252, 258, 323, 327, 352]. Suspension [318]. Switching [43]. Symbolic [149, 357]. Synchronization [8, 227, 275, 327]. Synchronizing [135]. Synchronous [350]. Synergism [168]. System [10, 16, 35, 57, 97, 103, 119, 131, 132, 139, 141– 145, 166, 180, 184, 224, 228, 247, 249, 295, 302, 311, 322, 359]. System/38 [142]. Systems [12, 17, 38–40, 42, 43, 52, 61, 76, 81, 86, 99, 114, 116, 118, 128, 145, 146, 168, 223, 225, 244, 248, 261, 263, 309, 310, 320, 351, 356]. Systolic [50, 157, 172, 174–181]. Systolic/Cellular [180].

Table [190]. Tables [266, 354]. talk [312]. target [85]. Task [19, 221, 292]. Task-Level [292]. Taulbee [258, 352]. Taxonomic [77]. Taxonomy [271]. Technique [91]. Techniques [64, 87, 120, 153, 177, 215]. Technology [83, 287, 345]. Temporal [151]. terrain [272]. terrains [316]. Test [300–302, 304]. Test-Generation [302]. Testability [6, 304]. Testbed [66–68]. Texas [193]. TEXnology [339]. Text [106]. Their [130]. TI [280]. Tightly [229]. Time [67, 78, 128, 230, 263, 320]. Titan [259]. Toast [143]. Tolerance [177]. Tolerant [96, 229]. Tools [10, 84, 87, 166, 335]. Top [104]. Top-Down [104]. Topology [329]. Trade [45, 277]. Trade-Offs [45, 277]. Traditional [250]. Transaction [229]. Transformations [319]. Translation [45]. Transport [333]. Turbulent [330]. Tutorial [239, 267]. Two [112]. Type [78].

U [70]. U-Interpreter [70]. Ubiquitous [326]. Underflow [27]. Union [260]. United [176]. Universal [361]. Unix [10, 28, 58]. Unix-Based [58]. unknown [316]. Update [200, 354]. USC [298]. use [5]. User [22, 125, 136, 284, 309]. Using [8, 44, 83, 112, 122, 234, 270, 317, 321].

Valued [238, 239, 241–243, 267]. variable [362]. VAX [45, 57]. VAX-11 [45]. VAX/VMS [57]. Vector [20, 251, 329, 341]. Vehicle [318]. Verification [107, 248, 252]. Very [49, 162]. Very-Large-Scale [49]. VHSIC [111]. via [362]. Videotex [21]. Views [136, 155]. Virtual [57]. Vision [69, 100, 183, 185, 322, 357, 359]. Visual [82, 123, 140, 235, 345]. Visualization [121, 328, 331, 333–338]. Visualizing [332, 346, 347]. Visually [33]. VLSI [1, 65, 84, 88, 108, 135, 233, 240, 265, 300, 307, 349]. VMS [57]. vs [225, 251].

Wang [312]. Wars [37]. Wavefront [173]. Were [60]. Whistle [301]. wide [12]. wide-band [12]. Windows [152]. Workbench [301]. Workstation [123]. world [82]. Write [247]. Write-Once [247]. Writing [127].

X [279, 341]. X-MP [279, 341]. Xerox [340].

Zeus [109]. ZS [324]. ZS-1 [324].

References

Foster:1980:DSV

[1] M. J. Foster and H. T. Kung. The design of special-purpose VLSI chips. Computer, 13(1):26–40, January 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coonen:1980:IGP

[2] Jerome T. Coonen. An implementation guide to a proposed standard for floating-point arithmetic. Computer, 13 (1):68–79, January 1980. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See errata in [24]. See [?, ?].

Thornton:1980:BNA

[3] J. E. Thornton. Back-end network approaches. Computer, 13(2):10–17, February 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chlamatac:1980:PIB

[4] I. Chlamatac, W. R. Franta, P. C. Patton, B., and Wells. Performance issues in back-end storage networks. Computer, 13(2):18–31, February 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Butland:1980:EGD

[5] J. Butland and S. D. Butland. An easyto-use graph drawing package. Computer, 13(2):69–80, February 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hayes:1980:TCM

[6] J. P. Hayes and E. J. McCluskey. Testability considerations in microprocessorbased design. Computer, 13(3):17– 26, March 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pradhan:1980:ECS

[7] D. K. Pradhan and J. J. Stiffler. Errorcorrecting codes and self-checking circuits. Computer, 13(3):27–37, March 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peacock:1980:DSU

[8] K. Peacock, E. Manning, and J. Wong. Synchronization of distributed simulation using broadcast algorithms. Computer, 13(4):3–10, April 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buzen:1980:MCQ

[9] Jeffrey P. Buzen and Peter J. Denning. Measuring and calculating queue length distributions. Computer, 13(4):

33–44, April 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1980:LDT

[10] Steven C. Johnson. Language development tools on the Unix system. Computer, 13(8):16–24, August 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leverett:1980:OPC

[11] Bruce W. Leverett, R. G. G. Cattell, S. O. Hobbs, J. M. Newcomer, A. H. Reiner, B. R. Schatz, and W. A. Wulf. An overview of the Production Quality Compiler-Compiler Project. Computer, 13(8):38–49, August 1980. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilkes:1980:IWL

[12] M. V. Wilkes. The impact of wideband local area communication systems on distributed computing. Computer, 13(9):22–25, September 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Isloor:1980:DPO

[13] S. S. Isloor and T. A. Marsland. The deadlock problem: An overview. Computer, 13(9):58–78, September 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flynn:1980:DIA

[14] M. Flynn. Directions and issues in architecture and language. Computer, 13(10): 5–22, October 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rodrigue:1980:PLS

[15] G. Rodrigue, E. D. Giroux, and M. Pratt. Perspectives on large-scale scientific computation. Computer, 13 (10):65–80, October 1980. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boberg:1980:PMS

[16] R. W. Boberg et al. Proposed Microcomputer System 796 Bus Standard. Computer, 13(10):89–105, October 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vick:1980:AAS

[17] C. R. Vick, S. P. Kartashev, and S. I. Kartashev. Adaptable architectures for super systems. Computer, 13(11):17– 35, November 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dennis:1980:DFS

[18] Jack B. Dennis. Data flow supercomputers. Computer, 13(11):48–56, November 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1980:TAD

[19] W. W. Chu, L. J. Holloway, M. Lan, and K. Efe. Task allocation in distributed data processing. Computer, 13(11):57– 69, November 1980. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kozdrowicki:1980:SGV

[20] E. W. Kozdrowicki and Douglas J. Theis. Second generation of vector supercomputers. Computer, 13(11):71– 83, November 1980. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Ball:1980:VN

[21] A. J. S. Ball, G. V. Bochmann, and J. Gecsei. Videotex networks. Computer, 13(12):8–14, December 1980. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1981:CGR

[22] W. Myers. Computer graphics: Reaching the user. Computer, 14(3):7–10, 14, 17, March 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stevenson:1981:PSB

[23] David Stevenson. A proposed standard for binary floating-point arithmetic. Computer, 14(3):51–62, March 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coonen:1981:EIG

[24] Jerome T. Coonen. Errata: An Implementation Guide to a Proposed Standard for Floating Point Arithmetic. Computer, 14(3):62, March 1981. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [2, ?, ?].

Cody:1981:APF

[25] William J. Cody, Jr. Analysis of proposals for the floating-point standard. Computer, 14(3):63–68, March 1981. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hough:1981:API

[26] David Hough. Applications of the proposed IEEE-754 standard for floating point arithmetic. Computer, 14(3):70– 74, March 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coonen:1981:UDN

[27] Jerome T. Coonen. Underflow and the denormalized numbers. Computer, 14 (3):75–87, March 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kernighan:1981:UPE

[28] B. W. Kernighan and J. R. Mashey. The Unix programming environment. Computer, 14(4):12–24, April 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Teitelman:1981:IPE

[29] Warren Teitelman and Larry Masinter. The Interlisp programming environment. Computer, 14(4):25–33, April 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zloof:1981:QOA

[30] M. M. Zloof. QBE/OBE: A language for office and business automation. Computer, 14(5):13–22, May 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stenning:1981:AEA

[31] V. Stenning, R. Froggatt, R. Gilbert, and E. Thomas. The Ada environment: A perspective. Computer, 14(6): 26–36, June 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sussman:1981:SLC

[32] Gerald Jay Sussman, Jack Holloway, Guy Lewis Steele, Jr., and Alan Bell. Scheme-79 — Lisp on a chip. Computer, 14(7):10–21, July 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Webb:1981:VIM

[33] Jon A. Webb and J. K. Aggarwal. Visually interpreting the motion of objects in space. Computer, 14(8):40–46, August 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohler:1981:FPA

[34] E. U. Cohler and J. E. Storer. Functionally parallel architecture for array processors. Computer, 14(9):28–36, September 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maron:1981:IAP

[35] N. Maron and T. A. Brengle. Integrating an array processor into a scientific computing system. Computer, 14(9):41– 44, September 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brode:1981:PFP

[36] B. Brode. Precompilation of Fortran programs to facilitate array processing. Computer, 14(9):46–51, September 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohen:1981:HWP

[37] Danny Cohen. On Holy Wars and a plea for peace. Computer, 14(10):48– 54, October 1981. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic). This is an entertaining account of the Big-Endian and Little-Endian problems of bit and byte ordering.

Pohm:1981:CMS

[38] A. V. Pohm and T. A. Smay. Computer memory systems. Computer, 14(10):93– 110, October 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:1981:PDB

[39] S. K. Chang and T. L. Kunii. Pictorial data base systems. Computer, 14(11): 13–21, November 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:1981:PQL

[40] N. S. Chang and K. S. Fu. Picture query languages for pictorial database systems. Computer, 14(11):23– 33, November 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zobrist:1981:PIP

[41] A. L. Zobrist and G. Nagy. Pictorial information processing of LANDSAT data for geographic analysis. Computer, 14 (11):34–41, November 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chock:1981:MDS

[42] Margaret Chock, Alfonso F. Cardenas, and Allen Klinger. Manipulating data structures in pictorial information systems. Computer, 14(11):43– 50, November 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dias:1981:PSI

[43] D. M. Dias and J. R. Jump. Packet switching interconnection networks for modular systems. Computer, 14(12):43– 53, December 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:1981:INU

[44] P.-Y. Chen, D. H. Lawrie, P.-C. Yew, and D. A. Padua. Interconnection networks using shuffles. Computer, 14(12): 55–63, December 1981. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Satyanarayanan:1981:DTV

[45] M. Satyanarayanan and D. P. Bhandarkar. Design trade-offs in VAX-11 translation buffer organization. Computer, 14(12):103–111, December 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Plattner:1981:MPE

[46] B. Plattner and J. Nievergelt. Monitoring program execution: A survey. Computer, 16(11):76–93, November 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lantz:1982:RIG

[47] K. A. Lantz, K. D. Gradischnig, J. A. Feldman, R. F., and Rashid. Rochester’s intelligent gateway. Computer, 11(10): 54–68, October 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haynes:1982:SHP

[48] L. S. Haynes, R. L. Lau, D. P. Siewiorek, and D. W. Mizel. A survey of highly parallel computing. Computer, 15(1):9– 24, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gottlieb:1982:NAV

[49] Allan Gottlieb and J. T. Schwartz. Networks and algorithms for very-large-scale parallel computation. Computer, 15(1):27–36, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kung:1982:WSA

[50] H. T. Kung. Why systolic architectures? Computer, 15(1):37–46, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmed:1982:HCC

[51] H. M. Ahmed, J. M. Delosme, and M. Morf. Highly concurrent computing structures for matrix arithmetic and signal processing. Computer, 15(1):65– 82, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Agerwala:1982:DS

[52] T. Agerwala and Arvind. DataFlow systems. Computer, 15(2):10–30, February 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ackerman:1982:DFL

[53] W. B. Ackerman. Data flow languages. Computer, 15(2):15–25, February 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davis:1982:DFP

[54] Alan L. Davis and Robert M. Keller. Data flow program graphs. Computer, 15(2):26–41, February 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Watson:1982:PDF

[55] I. Watson and J. Gurd. A practical data flow computer. Computer, 15(2): 51–57, February 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gajski:1982:SOD

[56] D. D. Gajski, D. A. Padua, D. J. Kuck, and R. H. Kuhn. A second opinion on data flow machines and languages. Computer, 15(2):58–69, February 1982. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Levy:1982:VMM

[57] H. M. Levy and P. H. Lipman. Virtual memory management in the VAX/ VMS operating system. Computer, 15 (3):35–41, March 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hwang:1982:ULC

[58] K. Hwang, W. J. Croft, G. H. Goble, B. W. Wah, F. A. Briggs, W. R. Simmons, and C. L. Coates. A Unix-Based local computer network with load balancing. Computer, 15(4):55–66, April 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kenville:1982:ODD

[59] Richard F. Kenville. Optical disk data storage. Computer, 15(7):21–25, July 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Copeland:1982:WIM

[60] George Copeland. What if mass storage were free? Computer, 15(7):27–35, July 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OLear:1982:SCM

[61] B. T. O’Lear and J. H. Choy. Software considerations in mass storage systems. Computer, 15(7):36–44, July 1982. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawrie:1982:EAF

[62] D. H. Lawrie, J. M. Randal, and R. R. Barton. Experiments with automatic file migration. Computer, 15(7):45–55, July 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shoch:1982:EEL

[63] John F. Shoch, Y. K. Dalal, David D. Redell, and Ronald C. Crane. Evolution of the Ethernet local computer network. Computer, 15(8):10–27, August 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boral:1982:ADF

[64] H. Boral and D. J. DeWitt. Applying data flow techniques to data base machines. Computer, 15(8):57– 63, August 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Reprinted in A. R. Hurson, L. L. Miller, and S. H. Pakzad, Parallel Architectures for Database Systems, IEEE Computer Society Press, Washington, DC, 1989.

Patterson:1982:VR

[65] D. A. Patterson and C. H. Sequin. A VLSI RISC. Computer, 15(9):8–21, September 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brayer:1982:TAD

[66] K. Brayer and V. Lafleur. A testbed approach to the design of a computer communication network. Computer, 15 (10):14–23, October 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McDonald:1982:FDT

[67] W. C. McDonald and R. W. Smith. A flexible distributed testbed for realtime applications. Computer, 15(10):25– 39, October 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gehringer:1982:CT

[68] E. F. Gehringer, A. K. Jones, and Z. Z. Segall. The Cm* testbed. Computer, 15 (10):40–49, October 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hall:1982:MCS

[69] Ernest L. Hall, James K. B. Tio, Charles A. McPherson, and Firooz A. Sadjadi. Measuring curved surfaces for robot vision. Computer, 15(12):42–54, December 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arvind:1982:U

[70] Arvind and Kim P. Gostelow. The U-Interpreter. Computer, 16(1):42– 49, January 1982. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fu:1983:ICA

[71] K. Hwang and K. Fu. Integrated computer architectures for image processing and database management. Computer, 16(1):51–60, January 1983. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atkinson:1983:APP

[72] M. P. Atkinson, P. J. Bailey, K. J. Chisholm, P. W. Cockshott, and R. Morrison. An approach to persistent programming. Computer, 16(4):360– 365, April 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Benhamou:1983:MIG

[73] Eric Benhamou and Judy Estrin. Multilevel internetworking gateways: Architecture and applications. Computer, 16 (9):27–34, September 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinden:1983:DII

[74] R. Hinden, J. Haverty, and A. Sheltzer. The DARPA Internet: Interconnecting heterogeneous computer networks with gateways. Computer, 16(9):38–48, September 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). papers box.

Thornton:1983:HNL

[75] James E. Thornton and Gary S. Christensen. Hyperchannel network links. Computer, 16(9):50–54, September 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Redell:1983:IEM

[76] David D. Redell and James E. White. Interconnecting electronic mail systems. Computer, 16(9):55–63, September 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brachman:1983:WII

[77] Ronald J. Brachman. What IS-A is and isn’t: An analysis of taxonomic links in semantic networks. Computer, 16 (10):30–36, October 1983. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schubert:1983:DTP

[78] Lenhart K. Schubert, Mary Angela Papalaskaris, and Jay Taugher. Determining type, part, color, and time relationships. Computer, 16(10):53– 60, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Special Issue on Knowledge Representation.

Brachman:1983:KAF

[79] Ronald J. Brachman, Richard E. Fikes, and Hector J. Levesque. KRYPTON: A functional approach to knowledge representation. Computer, 16(10):67– 73, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Special Issue on Knowledge Representation.

Davis:1983:RSB

[80] Randall Davis and Howard Shrobe. Representing structure and behavior of digital hardware. Computer, 16(10):75– 81, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mylopoulos:1983:BKS

[81] John Mylopoulos, Tetsutaro Shibahara, and John K. Tsotsos. Building knowledge-based systems: The PSN experience. Computer, 16(10):83–89, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Havens:1983:RKV

[82] W. Havens and A. Mackworth. Representing knowledge of the visual world. Computer, 16(10):90–96, October 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balzer:1983:STU

[83] Robert Balzer, Thomas E. Cheatham, Jr., and Cordell Green. Software technology in the 1990s: Using a new paradigm. Computer, 16(11):39–45, November 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gajski:1983:GEI

[84] D. Gajski and R. Kuhn. Guest Editors’ introduction: New VLSI tools. Computer, 16(12):11–14, December 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1984:BPS

[85] J. K. F. Lee and A. J. Smith. Branch prediction strategies and branch target buffer design. Computer, 17(1):6– 22, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gligor:1984:IHD

[86] V. D. Gligor and G. L. Luckenbaugh. Interconnecting heterogeneous database management systems. Computer, 17 (1):33–43, January 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boari:1984:MPT

[87] Maurelio Boari, Stefano Crespi-Reghizzi, Alberto Dapra, Francesco Maderna, and Antonio Natali. Multiplemicroprocessor programming techniques: MML, a new set of tools. Computer, 17(1):47–59, January 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grinberg:1984:CVA

[88] J. Grinberg, G. R. Nudd, and R. D. Etchells. A cellular VLSI architecture. Computer, 17(1):69–81, January 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weide:1984:PCI

[89] B. W. Weide, M. E. Brown, J. Ramanathan, and K. Schwan. Process control: Integration and design methodology support. Computer, 17(2):27– 32, February 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roussopoulos:1984:AMD

[90] N. Roussopoulos and R. T. Yeh. An adaptable methodology for database design. Computer, 17(5):64–80, May 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Welch:1984:THP

[91] Terry A. Welch. A technique for highperformance data compression. Computer, 17(6):8–19, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hsieh:1984:RSC

[92] W. N. Hsieh and I. Gitman. Routing strategies in computer networks. Computer, 17(6):46–56, June 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

MacDougall:1984:IPP

[93] Myron MacDougall. Instruction-level program and processor modeling. Computer, 17(7):14–24, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Snyder:1984:PPP

[94] L. Snyder. Parallel programming and the Poker programming environment. Computer, 17(7):27–36, July 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keller:1984:SPR

[95] R. M. Keller and F. C. H. Lin. Simulated performance of a reduction-based multiprocessor. Computer, 17(7):70–82, July 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sarrazin:1984:FSM

[96] D. B. Sarrazin and M. Malek. Faulttolerant semiconductor memories. Computer, 17(8):49–56, August 1984. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mittal:1984:PAK

[97] S. Mittal, B. Chandrasekaran, and J. Sticklen. Patrec: A knowledgedirected database for a diagnostic expert system. Computer, 17(9):51–58, September 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

MacHover:1984:ICG

[98] C. MacHover and W. Myers. Interactive computer graphics. Computer, 17(10): 145–161, October 1984. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1984:AOS

[99] R. L. Brown and P. J. Denning. Advanced operating systems. Computer, 17(10):173–190, October 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fu:1984:PRC

[100] K. S. Fu and A. Rosenfeld. Pattern recognition and computer vision. Computer, 17(10):274–282, October 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Glinert:1984:PIG

[101] E. P. Glinert and S. L. Tanimoto. PICT: An interactive graphical programming environment. Computer, 17(11):7–25, November 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shemer:1984:GDC

[102] J. Shemer and P. M. Neches. The genesis of a database computer. Computer, 17(11):42–??, November 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laurent:1984:GAG

[103] D. Laurent and S. Motet. GEOMATIC:

A 3-D graphic relief simulation system. Computer, 17(12):25–30, December 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yoshiura:1984:TCM

[104] H. Yoshiura, K. Fujimura, and T. L.

Kunii. Top-down construction of 3-D mechanical object shapes from engineering drawings. Computer, 17(12):32– 40, December 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:1984:ISM

[105] T. C. Woo. Interfacing solid modeling to CAD and CAM: Data structures and algorithms for decomposing a solid. Computer, 17(12):44–49, December 1984. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1985:CKT

[106] Y. Chu. Chinese/Kanji text and data processing. Computer, 18(1):10– 66, January 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maruyama:1985:HV

[107] F. Maruyama and M. Fujita. Hardware verification. Computer, 18(2): 22–32, February 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). U. S. Copyright Clearance Center Code: 0018-9162/85/0200-0022$01.00.

Suzuki:1985:CPE

[108] N. Suzuki. Concurrent PROLOG as an efficient VLSI design language. Computer, 18(2):33–40, February 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). U. S. Copyright Clearance Center Code: 0018-9162/85/0200-0033$01.00.

Lieberherr:1985:ZAL

[109] K. J. Lieberherr and S. M. German. Zeus: A language for expressing algorithms in hardware. Computer, 18(2): 55–65, February 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Zeus.

Piloty:1985:CPC

[110] R. Piloty and D. Borrione. The CON-LAN project: concepts, implementations and applications. Computer, 18 (2):81–92, February 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shahdad:1985:VHD

[111] M. Shahdad, R. Lipsett, F. E. Marschner, K. Sheehan, H. Cohen, R. Waxman, and D. Ackley. VHSIC hardware description language. Computer, 18(2):94–103, February 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). VHDL.

Berman:1985:CRT

[112] S. Berman, P. Parikh, and C. S. G. Lee. Computer recognition of two overlapping parts using a single camera. Computer, 18(3):70–80, March 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1985:NCA

[113] Anonymous. New configuration added, price cut, for processor. Computer, 18 (5):122–??, May 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gajski:1985:EIM

[114] Daniel D. Gajski and Jih-Kwon Peir. Essential issues in multiprocessor systems. Computer, 18(6):9–27, June 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneck:1985:PPP

[115] P. B. Schneck, D. Austin, S. L. Squires, J. Lehmann, D. Mizell, and K. Wallgren. Parallel processor programs in the Federal Government. Computer, 18(6): 43–56, June 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murakami:1985:RPM

[116] K. Murakami, T. Kakuta, R. Onai, and N. Ito. Research on parallel machine architecture for fifth-generation computer systems. Computer, 18(6): 76–92, June 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Month June.

Wah:1985:MCS

[117] Benjamin W. Wah, G. J. Li, and C. F.

Yu. Multiprocessing of combinatorial search problems. Computer, 18(6):93– 108, June 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:1985:TRM

[118] S. W. Miller and M. W. Collins. Toward a reference model of mass storage systems. Computer, 18(7):9–22, July 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ammon:1985:HLJ

[119] G. Ammon, J. Calabria, and D. Thomas. A high-speed, large-capacity, jukebox optical disk system. Computer, 18(7): 36–46, July 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raeder:1985:SCG

[120] G. Raeder. A survey of current graphical programming techniques. Computer, 18(8):11–25, August 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1985:PVG

[121] G. P. Brown, R. T. Carling, C. F. Herot, D. A. Kramlich, and P. Souza. Program visualization: Graphical support for software development. Computer, 18 (8):27–35, August 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

London:1985:APU

[122] R. L. London and R. A. Duisberg. Animating programs using Smalltalk. Computer, 18(8):61–71, August 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Melamed:1985:VSP

[123] B. Melamed and R. J. T. Morris. Visual simulation: The performance analysis workstation. Computer, 18(8):87– 94, August 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Colwell:1985:ISB

[124] Robert P. Colwell, Charles Y. Hitchcock, III, E. Douglas Jensen, H. M. Brinkley Sprunt, and Charles P. Kollar. Instruction sets and beyond: Computers, complexity, and controversy. Computer, 18 (9):8–19, September 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coutaz:1985:AUI

[125] Joëlle Coutaz. Abstractions for user interface design. Computer, 18(9):21– 34, September 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hartzband:1985:EKR

[126] D. J. Hartzband and F. J. Maryanski.

Enhancing knowledge representation in engineering databases. Computer, 18(9): 39–??, September 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yankelovich:1985:RWE

[127] Nicole Yankelovich, Norman Meyrowitz, and Andries van Dam. Reading and writing the electronic book. Computer, 18(10):15–30, October 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sarin:1985:CRC

[128] S. Sarin and I. Greif. Computer-based real-time conferencing systems. Computer, 18(10):33–45, October 1985. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horak:1985:ODA

[129] Wolfgang Horak. Office Document Architecture and Office Document Interchange formats: Current status of international standardization. Computer, 18 (10):50–60, October 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bono:1985:SGS

[130] Peter R. Bono. A survey of graphics standards and their role in information interchange. Computer, 18(10):63– 75, October 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reynolds:1985:DEM

[131] J. K. Reynolds, J. B. Postel, A. R. Katz, G. G. Finn, and A. L. DeSchon. The DARPA experimental multimedia mail system. Computer, 18(10):82– 91, October 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sakata:1985:DIM

[132] S. Sakata and T. Ueda. A distributed interoffice mail system. Computer, 18(10): 106–116, October 1985. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carlson:1985:APD

[133] W. W. Carlson and K. Hwang. Algorithmic performance of dataflow multiprocessors. Computer, 18(12):30–40, December 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yalamanchili:1985:RSP

[134] S. Yalamanchili and J. K. Aggarwal. Reconfiguration strategies for parallel architectures. Computer, 18(12):44–61, December 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fisher:1985:SLV

[135] A. L. Fisher and H. T. Kung. Synchronizing large VLSI processor arrays. Computer, C-34(8):734–740, August 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Navathe:1986:IUV

[136] S. Navathe, R. Elmasri, and J. Larson.

Integrating user views in database design. Computer, 19(1):50–62, January 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goguen:1986:RIS

[137] Joseph Goguen. Reusing and interconnecting software components. Computer, 19(2):16–28, February 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Reprinted in Tutorial: Software Reusability, Peter Freeman, editor, IEEE Computer Society, 1987, pages 251–263, and in Domain Analysis and Software Systems Modelling, Ruben Prieto-Diaz and Guillermo Arango, editors, IEEE Computer Society, 1991, pages 125–137.

Agrawal:1986:EPM

[138] D. P. Agrawal, V. K. Janakiram, and G. C. Pathak. Evaluating the performance of multicomputer configurations.

Computer, 19(5):23–37, May 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fisher:1986:IDD

[139] D. A. Fisher and R. M. Weatherly. Issues in the design of a distributed operating system for Ada. Computer, 19(5): 38–47, May 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Larson:1986:VAB

[140] James A. Larson. A visual approach to browsing in a database environment. Computer, 19(6):62–71, June 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bruno:1986:RSS

[141] G. Bruno, A. Elia, and P. Laface. A rulebased system to schedule production. Computer, 19(7):32–??, July 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stroebel:1986:CPE

[142] G. J. Stroebel, R. D. Baxter, and M. J. Denney. A capacity planning expert system for IBM System/38. Computer, 19 (7):42–??, July 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Talukdar:1986:TPS

[143] S. N. Talukdar, E. Cardozo, and L. V. Leao. Toast: The power system operator’s assistant. Computer, 19(7): 53–??, July 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1986:KCM

[144] David C. Brown and B. Chandrasekaran. Knowledge and control for a mechanical design expert system. Computer, 19(7):92–100, July 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mittal:1986:PES

[145] S. Mittal, C. L. Dym, and M. Morjaria.

Pride: An expert system for the design of paper handling systems. Computer, 19(7):102–??, July 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chien:1986:ESS

[146] Y.-T. Chien and J. Liebowitz. Expert systems in the SDI environment. Computer, 19(7):115–??, July 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mundie:1986:PPA

[147] D. A. Mundie and D. A. Fisher. Parallel processing in Ada. Computer, 19(8): 20–25, August 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahuja:1986:DPL

[148] Sudhir Ahuja, Nicholas Carriero, and David Gelernter. Domesticating parallelism: Linda and friends. Computer, 19 (8):26–34, August 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Halstead:1986:PSC

[149] R. H. Halstead. Parallel symbolic computing. Computer, 19(8):35–43, August 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hudak:1986:PP

[150] P. R. Hudak. Para-functional programming. Computer, 19(8):60–70, August 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Snodgrass:1986:TD

[151] R. T. Snodgrass and I. Ahn. Temporal databases. Computer, 19(9):35–42, September 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1986:CEI

[152] B. A. Myers. A complete and efficient implementation of covered windows. Computer, 19(9):57–67, September 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roussopoulos:1986:PTD

[153] N. Roussopoulos and H. Kang. Principles and techniques in the design of ADMS. Computer, 19(12):19–25, December 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mark:1986:MM

[154] L. Mark and N. Roussopoulos. Metadata management. Computer, 19(12): 26–36, December 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wiederhold:1986:VOD

[155] G. Wiederhold. Views, objects, and databases. Computer, 19(12):37–44, December 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Christodoulakis:1986:DPC

[156] S. Christodoulakis and C. Faloutsos. Design and performance considerations for an optical disk-based, multimedia object server. Computer, 19(12):45–56, December 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moldovan:1986:PMA

[157] D. I. Moldovan and J. A. B. Fortes. Partitioning and mapping algorithms into fixed-size systolic arrays. Computer, C-35(1):1–12, January 1986. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hwang:1987:CAA

[158] K. Hwang, J. Ghosh, and R. Chowkwanyun. Computer architectures for artificial intelligence processing. Computer, 20(1):19–27, January 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ungar:1987:WPS

[159] D. Ungar and D. Patterson. What price Smalltalk? Computer, 20(1):67– 74, January 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fahlman:1987:CAA

[160] S. E. Fahlman and Geoffrey E. Hinton. Connectionist architectures for artificial intelligence. Computer, 20(1):100– 109, January 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ketabchi:1987:MMC

[161] M. A. Ketabchi and V. Berzins. Modeling and managing CAD databases. Computer, 20(2):93–102, February 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berra:1987:CAS

[162] P. B. Berra, Soon M. Chung, and Nabil I. Hachem. Computer architecture for a surrogate file to a very large data/knowledge base. Computer, 20(3): 25–32, March 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pleszkun:1987:ALM

[163] A. R. Pleszkun and M. J. Thazhuthaveetil. The architecture of LISP machines. Computer, 20(3):35–44, March 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dongarra:1986:HDM

[164] J. Dongarra. How do the ‘minisupers’ stack up? Computer, 19 (3):93, 100, March 1986. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://www.netlib.org/utk/people/
                                        JackDongarra/PAPERS/How-Do-the-
                                        Minisupers-Stack-Up.pdf.

Brooks:1987:NSB

[165] Frederick P. Brooks, Jr. No silver bullet: essence and accidents of software engineering. Computer, 20(4):10–19, April 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gevarter:1987:NEC

[166] W. B. Gevarter. The nature and evaluation of commercial expert system building tools. Computer, 20(5):24–41, May 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karp:1987:PP

[167] Alan H. Karp. Programming for parallelism. Computer, 20(5):43–57, May 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milutinovic:1987:ACS

[168] V. Milutinovic, D. Fura, W. Helbig, and J. Linn. Architecture/compiler synergism in GaAs computer systems. Computer, 20(5):72–93, May 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhuyan:1987:GEI

[169] L. N. Bhuyan. Guest Editor’s introduction: Interconnection networks for parallel and distributed processing. Computer, 20(6):9–13, June 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:1987:ASM

[170] Vijay P. Kumar and Sudhakar M.

Reddy. Augmented shuffle-exchange multistage interconnection networks. Computer, 20(6):30–40, June 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Varma:1987:OCN

[171] A. Varma, A. A. Sawchuk, C. S.

Raghavendra, and B. K. Jenkins. Optical crossbar networks. Computer, 20(6): 50–60, June 1987. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Fortes:1987:SAC

[172] J. A. B. Fortes and B. W. Wah. Systolic arrays — from concept to implementation. Computer, 20(7):12–17, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kung:1987:WAP

[173] S. Y. Kung, S. N. Lo, and J. N. Hwang. Wavefront array processors — concept to implementation. Computer, 20(7): 18–34, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foulser:1987:SMA

[174] D. E. Foulser and R. Schreiber. The Saxpy Matrix-1: A general-purpose systolic computer. Computer, 20(7):35– 44, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drake:1987:SAS

[175] B. L. Drake, F. T. Luk, J. M. Speiser, and J. J. Symanski. SLAPP: A systolic linear algebra parallel processor. Computer, 20(7):45–50, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McCanny:1987:SSA

[176] J. V. McCanny and J. G. McWhirter. Some systolic array developments in the united kingdom. Computer, 20(7): 51–64, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abraham:1987:FTT

[177] J. A. Abraham, P. Banerjee, C. Chen, W. K. Fuchs, S. Y. Kuo, and A. L. N. Reddy. Fault tolerance techniques for systolic arrays. Computer, 20(7):65– 76, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Navarro:1987:PES

[178] J. J. Navarro, J. M. Llaberia, and M. Valero. Partitioning: An essential step in mapping algorithms into systolic array processors. Computer, 20(7): 77–90, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hein:1987:DGS

[179] C. E. Hein, R. M. Zieger, and J. A. Urbano. The design of a GaAs systolic array for an adaptive null steering beamforming controller. Computer, 20(7): 92–93, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nash:1987:SCS

[180] J. G. Nash, K. W. Przytula, and S. Hansen. The systolic/cellular system for signal processing. Computer, 20(7): 96–97, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:1987:ISA

[181] W. T. Lin, C. Y. Chin, and C. Y. Ho.

Integrating systolic arrays into a supersystem. Computer, 20(7):100–101, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koren:1987:CID

[182] Israel Koren. The concept and implementation of data-driven processor arrays. Computer, 20(7):102–103, July 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhanu:1987:COR

[183] Bir Bhanu and Chih-Cheng Ho. CADbased 3D object representation for robot vision. Computer, 20(8):19–35, August 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roach:1987:CSB

[184] John W. Roach, Praveen K. Paripati, and Jeffrey S. Wright. A CAD system based on spherical dual representations. Computer, 20(8):37–44, August 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shneier:1987:PBV

[185] Michael O. Shneier, Ronald Lumia, and Martin Herman. Prediction-based vision for robot control. Computer, 20(8):46– 55, August 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horaud:1987:MSP

[186] Radu Horaud and Thomas Skordas. Model-based strategy planning for recognizing partially occluded parts. Computer, 20(8):58–65, August 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gunnarsson:1987:CML

[187] Kristjan T. Gunnarsson and Friedrich B. Prinz. CAD model-based localization of parts in manufacturing. Computer, 20 (8):66–74, August 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stevenson:1987:MPC

[188] Charles N. Stevenson. Model-based programming and control of robot manipulators. Computer, 20(8):76–84, August 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:FC

[189] Anonymous. [front cover]. Computer, 20(9):c1–c2, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:TC

[190] Anonymous. Table of contents. Computer, 20(9):2–3, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Russo:1987:CSP

[191] R. Russo and M. Smith. Computer Society President’s message. Computer, 20 (9):6–7, September 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shriver:1987:ECM

[192] B. D. Shriver. Editor-in-Chief’s message. Computer, 20(9):10, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:FJC

[193] Anonymous. The 1987 Fall Joint Computer Conference Infomart — Dallas, Texas, October 25–29, 1987. Computer, 20(9):12–15, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conklin:1987:HIS

[194] Jeff Conklin. Hypertext: An introduction and survey. Computer, 20(9):17– 41, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boehm:1987:ISP

[195] Boehm. Improving software productivity. Computer, 20(9):43–57, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gimarc:1987:SRP

[196] C. E. Gimarc and V. M. Milutinovic. A survey of RISC processors and computers of the mid-1980s. Computer, 20(9): 59–69, September 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flynn:1987:NCM

[197] M. J. Flynn, C. L. Mitchell, and J. M. Mulder. And now a case for more complex instruction sets. Computer, 20(9): 71–83, September 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoare:1987:OSF

[198] C. A. R. Hoare. An overview of some formal methods for program design. Computer, 20(9):85–91, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:S

[199] Anonymous. Standards. Computer, 20 (9):94–97, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:U

[200] Anonymous. Update. Computer, 20(9):

98–99, September 1987. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:CSN

[201] Anonymous. Computer Society news.

Computer, 20(9):99–102, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:NPR

[202] Anonymous. New product reviews.

Computer, 20(9):103–106, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:NP

[203] Anonymous. New products. Computer, 20(9):107–113, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:IA

[204] Anonymous. IC announcements. Computer, 20(9):114, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:MA

[205] Anonymous. Microsystem announcements. Computer, 20(9):115, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:IIC

[206] Anonymous. IEEE International Conference on Computer-Aided Design. Computer, 20(9):116–117, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:Ca

[207] Anonymous. Conferences. Computer, 20 (9):118–119, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:Cb

[208] Anonymous. Calendar. Computer, 20 (9):119–120, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:CP

[209] Anonymous. Call for papers. Computer, 20(9):121–122, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:CO

[210] Anonymous. Career opportunities. Computer, 20(9):122–123, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:BR

[211] Anonymous. Book reviews. Computer, 20(9):124–126, September 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:NL

[212] Anonymous. New literature. Computer, 20(9):127, September 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1987:CSM

[213] Anonymous. The Computer Society: a member society of the Institute of Electrical and Electronics Engineers, Inc. Computer, 20(9):129, September 1987.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1987:PQL

[214] Harold S. Stone. Parallel querying of large databases: A case study. Computer, 20(10):11–21, October 1987. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berra:1987:OTD

[215] P. B. Berra and N. B. Troullinos.

Optical techniques and data/knowledge base machines. Computer, 20(10):59– 70, October 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Henderson:1987:IDP

[216] P. B. Henderson and D. Notkin. Integrated design and programming environments. Computer, 20(11):12–16, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dart:1987:SDE

[217] S. A. Dart, R. J. Ellison, P. H. Feiler, and N. A. Habermann. Software development environments. Computer, 20 (11):18–28, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reps:1987:LPP

[218] T. Reps and T. Teitelbaum. Language processing in program editors. Computer, 20(11):29–40, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goguen:1987:FPE

[219] Joseph Goguen and Mark Moriconi. Formalization in programming environments. Computer, 20(11):55–64, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carle:1987:PES

[220] Alan Carle, Keith D. Cooper, Robert T. Hood, Ken Kennedy, Linda Torczon, and Scott K. Warren. A practical environment for scientific programming. Computer, 20(11):75–89, November 1987. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Denning:1988:CDP

[221] P. J. Denning, D. E. Comer, D. Gries, M. C. Mulder, A. Tucker, A. J. Turner, and P. R. Young. Computing as a discipline: Preliminary report of the ACM Task Force on the Core of Computer Science. Computer, 20(1):41, February 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also final report [288].

Chen:1988:MRD

[222] P. Chen and M. A. Harrison. Multiple representation document development. Computer, 21(1):15–31, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1988:SDE

[223] J. Johnson and R. J. Beach. Styles in document editing systems. Computer, 21(1):32–47, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schlichter:1988:FAP

[224] J. H. Schlichter and L. J. Miller. FolioPub: A publication management system. Computer, 21(1):61–69, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marchionini:1988:FFV

[225] G. Marchionini and B. Shneiderman.

Finding facts vs. browsing knowledge in hypertext systems. Computer, 21(1):70– 80, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yankelovich:1988:ICC

[226] Nicole Yankelovich, Bernard Hann, Norman Meyrowitz, and S. M. Drucker. Intermedia: The concept and the construction of a seamless information environment. Computer, 21(1):81–97, January 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dubois:1988:SCE

[227] Michel Dubois, Christoph Scheurich, and Fayé A. Briggs. Synchronization, coherence, and event ordering in multiprocessors. Computer, 21(2):9–21, February 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ousterhout:1988:SNO

[228] John K. Ousterhout, A. R. Cherenson, Fred Douglis, Michael N. Nelson, and Brent B. Welch. The Sprite network operating system. Computer, 21(2):23– 36, February 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bernstein:1988:SAF

[229] P. A. Bernstein. Sequoia: A faulttolerant tightly coupled multiprocessor for transaction processing. Computer, 21 (2):37–45, February 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lang:1988:OPA

[230] Gordon R. Lang, Moez Dharssi, Fred M. Longstaff, Philip S. Longstaff, Peter A. S. Metford, and Malcolm T. Rimmer. An optimum parallel architecture for high-speed real-time digital signal processing. Computer, 21(2):47– 57, February 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moffett:1988:SAC

[231] Jonathan D. Moffett and Morris S. Sloman. The source of authority for commercial access control. Computer, 21 (2):59–69, February 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Widrow:1988:NNA

[232] B. Widrow and R. Winter. Neural nets for adaptive filtering and adaptive pattern recognition. Computer, 21(3):25– 39, March 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Graf:1988:VIN

[233] Hans P. Graf, Lawrence D. Jackel, and Wayne E. Hubbard. VLSI implementation of a neural network model. Computer, 21(3):41–49, March 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hutchinson:1988:CMU

[234] James Hutchinson, Christof Koch, Jin Luo, and Carver Mead. Computing motion using analog and binary resistive networks. Computer, 21(3):52– 63, March 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fukushima:1988:NNF

[235] Kunihiko Fukushima. Neural network for visual pattern recognition. Computer, 21(3):65–75, March 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carpenter:1988:AAP

[236] Gail A. Carpenter and Stephen Grossberg. The ART of adaptive pattern recognition by a self-organizing neural network. Computer, 21(3):77–88, March 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feldmann:1988:CSN

[237] Jerome A. Feldmann, Mark A. Fanty, and Nigel H. Goddard. Computing with structured neural networks. Computer, 21(3):91–103, March 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Butler:1988:MLG

[238] J. T. Butler. Multiple-valued logic:

Guest Editor’s introduction and bibliography. Computer, 21(4):13–??, April 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1988:MLA

[239] K. C. Smith. Multiple-valued logic: A tutorial and appreciation. Computer, 21(4):17–27, April 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Etiemble:1988:CBM

[240] D. Etiemble and M. Israel. Comparison of binary and multivalued ICs according to VLSI criteria. Computer, 21(4): 28–42, April 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kameyama:1988:MCM

[241] M. Kameyama, S. Kawahito, and T. Higuchi. A multiplier chip with multiple-valued bidirectional currentmode logic circuits. Computer, 21(4): 43–57, April 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Butler:1988:MCC

[242] J. T. Butler and H. G. Kerkhoff. Multiple-valued CCD circuits. Computer, 21(4):58–70, April 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sasao:1988:MLO

[243] T. Sasao. Multiple-valued logic and optimization of programmable logic arrays. Computer, 21(4):71–82, April 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nichols:1988:MMS

[244] Kathleen M. Nichols and John T. Edmark. Modeling multicomputer systems with PARET. Computer, 21(5): 39–48, May 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bershad:1988:RCF

[245] Brian N. Bershad and Henry M. Levy. A remote computation facility for a heterogeneous environment. Computer, 21(5): 50–60, May 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boehm:1988:SMS

[246] Barry W. Boehm. A spiral model of software development and enhancement. Computer, 21(5):61–72, May 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gait:1988:OFC

[247] Jason Gait. The optical file cabinet:

A random-access file system for writeonce optical disks. Computer, 21(6): 11–22, June 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mills:1988:SRV

[248] H. D. Mills. Stepwise refinement and verification in box-structured systems. Computer, 21(6):23–37, June 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rayfield:1988:SAS

[249] J. T. Rayfield and H. F. Silverman. System and application software for the Armstrong multiprocessor. Computer, 21(6):38–52, June 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Potter:1988:TSH

[250] Walter D. Potter and Robert P. Trueblood. Traditional, semantic, and hypersemantic approaches to data modeling.

Computer, 21(6):53–63, June 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:1988:PVF

[251] Bowen Liu and Nelson Strother. Programming in VS Fortran on the IBM 3090 for maximum vector performance. Computer, 21(6):65–77, June 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Camurati:1988:FVH

[252] P. Camurati and P. Prinetto. Formal verification of hardware correctness: Introduction and survey of current research. Computer, 21(7):8–20, July 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hooton:1988:EPD

[253] A. R. Hooton, U. Agueero, and S. Dasgupta. An exercise in plausibility-driven design. Computer, 21(7):21–33, July 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Steenkiste:1988:LRP

[254] P. Steenkiste and J. Hennessy. Lisp on a reduced-introduction-set processor: Characterization and optimization. Computer, 21(7):34–46, July 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athas:1988:MMP

[255] William C. Athas and Charles L. Seitz. Multicomputers: Message-passing concurrent computers. Computer, 21(8):9– 24, August 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tucker:1988:AAC

[256] Lewis W. Tucker and George G. Robertson. Architecture and applications of the Connection Machine. Computer, 21(8): 26–38, August 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:1988:APM

[257] Charles Rich and Richard C. Waters. Automatic programming: Myths and prospects. Computer, 21(8):40– 51, August 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gries:1988:TS

[258] David Gries and Dorothy Marsh. The 1986–1987 Taulbee Survey. Computer, 21(8):53–63, August 1988. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Diede:1988:TGS

[259] Tom Diede, Carl F. Hagenmaier, Glen S.

Miranker, Jonathan J. Rubinstein, and William S. Worley, Jr. The Titan Graphics Supercomputer architecture. Computer, 21(9):13–28, 30, September 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolcott:1988:HCS

[260] Peter Wolcott and Seymour E. Goodman. High-speed computers of the Soviet Union. Computer, 21(9):32–41, September 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leung:1988:FCE

[261] K. S. Leung and W. Lam. Fuzzy concepts in expert systems. Computer, 21 (9):43–56, September 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1988:NCS

[262] W. Myers and M. Potes. National Conference on Strategic Management of R&D. Computer, 21(9):58–67, September 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stankovic:1988:MAR

[263] J. A. Stankovic. Misconceptions about real-time computing — A serious problem for next-generation systems. Computer, 21(10):10–19, October 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Comment 1 by schlenk, Mon Dec 12 13:13:04 1988 An interesting outlook on real-time systems that are integrated into a programming environment is given. The author shows why ADA tasks are not sufficient (a high priority task has to wait for a lower priority task) and suggests an integrated approach: Time management on the language level, schedulability check and support for distributed programs.

Doran:1988:AMA

[264] R. W. Doran. Amdahl multiple-domain architecture. Computer, 21(10):20– 28, October 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koren:1988:DVA

[265] Israel Koren, Bilha Mendelson, Irit Peled, and Gabriel M. Silberman. A data-driven VLSI array for arbitrary algorithms. Computer, 21(10):30–43, October 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1988:HDS

[266] Ted G. Lewis and Curtis R. Cook. Hashing for dynamic and static internal tables. Computer, 21(10):45–56, October 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). The authors survey the classical hashing function approach to information retrieval and show how general hashing techniques exchange speed for memory. It is a tutorial paper that covers, among other topics, dynamic and static hash tables, perfect hashing, and minimal perfect hashing.

Lie:1988:AML

[267] Y.-C. D. Lie. Appendum to ‘Multiplevalued logic: a tutorial and appreciation’ (April 17–27, 1988). Computer, 21(11): 9, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:1988:PAR

[268] C. Rich and R. C. Waters. The Programmer’s Apprentice: a research overview. Computer, 21(11):10–25, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stenstrom:1988:RCS

[269] P. Stenström. Reducing contention in shared-memory multiprocessors. Computer, 21(11):26–37, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Britcher:1988:UII

[270] R. N. Britcher. Using inspections to investigate program correctness. Computer, 21(11):38–44, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Skillicorn:1988:TCA

[271] David B. Skillicorn. A taxonomy for computer architectures. Computer, 21 (11):46–57, November 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Webster:1988:MDI

[272] D. E. Webster. Mapping the design information representation terrain. Computer, 21(12):8–23, December 1988. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hill:1988:CDC

[273] M. D. Hill. A case for direct-mapped caches. Computer, 21(12):25–40, December 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaffner:1988:SSD

[274] Stuart C. Schaffner and Martha Borkan. Segue: Support for distributed graphical interfaces. Computer, 21(12):42– 55, December 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Detlefs:1988:ISR

[275] David L. Detlefs, Maurice P. Herlihy, and Jeannette M. Wing. Inheritance of synchronization and recovery properties in Avalon/C++. Computer, 21(12):57– 69, December 1988. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1988:SS

[276] Anonymous. Software for supercomputers. Computer, 21(12):70–74, December 1988. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rau:1989:CDS

[277] B. Ramakrishna Rau, D. W. L. Yen, Wei Yen, and R. A. Towle. The Cydra 5 departmental supercomputer: design philosophies, decisions, and tradeoffs. Computer, 22(1):12–35, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1989:EDC

[278] T. Jones. Engineering design of the Convex C2. Computer, 22(1):36–44, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

August:1989:CXB

[279] M. C. August, G. M. Brost, C. C. Hsiung, and A. J. Schiffleger. Cray X-MP: The birth of a supercomputer. Computer, 22(1):45–52, January 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cragon:1989:TAS

[280] H. G. Cragon and W. J. Watson. The TI Advanced Scientific Computer. Computer, 22(1):55–64, January 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maytal:1989:DCG

[281] B. Maytal, S. Iacobovici, D. B. Alpert, D. Biran, J. Levy, and S. Y. Tov. Design considerations for a general-purpose microprocessor. Computer, 22(1):66– 76, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1989:PA

[282] R. B. Lee. Precision architecture. Computer, 22(1):78–91, January 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1989:SPS

[283] W. Myers. Software pivotal to strategic defense. Computer, 22(1):92–97, January 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Linton:1989:CUI

[284] Mark A. Linton, John M. Vissides, and Paul R. Calder. Composing user interfaces with InterViews. Computer, 22 (2):8–22, February 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). From Computing Reviews: “InterViews objects are combined through a model similar to the boxes and glues in Knuth’s TEX, which allows interface objects to be combined with little concern for undesired interactions.”.

Bhuyan:1989:PMI

[285] L. N. Bhuyan, Qing Yang, and D. P. Agrawal. Performance of multiprocessor interconnection networks. Computer, 22 (2):25–37, February 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Richardson:1989:COA

[286] Stephen Richardson and Mahadevan Ganapathi. Code optimization across procedures. Computer, 22(2):42–50, February 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ganapathi:1989:IAC

[287] M. Ganapathi and G. O. Mendal. Issues in Ada compiler technology. Computer, 22(2):52–60, February 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Denning:1989:CD

[288] P. J. Denning, D. E. Comer, D. Gries, M. C. Mulder, A. Tucker, A. J. Turner, and P. R. Young. Computing as a discipline. Computer, 22(2):63–70, February 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also preliminary report [221].

Gries:1989:SRD

[289] D. Gries, T. Walker, and P. R. Young.

1988 Snowbird report: a discipline matures (computer science education). Computer, 22(2):72–75, February 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilfong:1989:RA

[290] G. T. Wilfong. Robotics and automation. Computer, 22(3):6–8, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharir:1989:AMP

[291] M. Sharir. Algorithmic motion planning in robotics. Computer, 22(3):9– 19, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lozano-Perez:1989:TPP

[292] T. Lozano-Perez, J. L. Jones, E. Mazer, and P. A. O’Donnell. Task-level planning of pick-and-place robot motions. Computer, 22(3):21–29, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoffmann:1989:PAR

[293] C. M. Hoffmann. The problems of accuracy and robustness in geometric computation. Computer, 22(3):31–39, 41, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cox:1989:EHR

[294] I. J. Cox and N. H. Gehani. Exception handling in robotics. Computer, 22(3): 43–49, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Allen:1989:CRR

[295] P. Allen, P. Michelman, and K. Roberts. Current research in robotics and automation — an intelligent grasping system. Computer, 22(3):50–52, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eicker:1989:CRR

[296] P. J. Eicker and D. R. Strip. Current research in robotics and automation — automated planning and programming for robotic batch mechanical assembly. Computer, 22(3):53–54, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hopcroft:1989:CRR

[297] J. E. Hopcroft. Current research in robotics and automation — electronic prototyping. Computer, 22(3): 55–57, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bekey:1989:CRR

[298] G. A. Bekey. Current research in robotics and automation-robotics research at USC. Computer, 22(3):58– 60, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keefe:1989:SQS

[299] T. F. Keefe, M. B. Thuraisingham, and W. T. Tsai. Secure queryprocessing strategies. Computer, 22(3): 63–70, March 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhattacharya:1989:HTG

[300] Debashis Bhattacharya, Brian T. Murray, and John P. Hayes. High-level test generation for VLSI. Computer, 22(4): 16–24, April 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Renous:1989:WWT

[301] R. Renous, G. M. Silberman, and I. Spillinger. Whistle: A workbench for test development of library-based designs. Computer, 22(4):27–41, April 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheng:1989:GAT

[302] W.-T. Cheng and T. J. Chakraborty. Gentest: An automatic test-generation system for sequential circuits. Computer, 22(4):43–49, April 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Heap:1989:GSC

[303] M. A. Heap and W. A. Rogers. Generating single-stuck-fault coverage from a collapsed-fault set. Computer, 22(4): 51–57, April 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hallenbeck:1989:TEA

[304] J. J. Hallenbeck, J. R. Cybrynski, N. Kanopoulos, T. Markas, and N. Vasanthavada. The Test Engineer’s Assistant: A support environment for hardware design for testability. Computer, 22(4):59–68, April 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tanik:1989:RPS

[305] M. M. Tanik and R. T. Yeh. Rapid prototyping in software development. Computer, 22(5):9–12, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Luqi:1989:SET

[306] Luqi. Software evolution through rapid prototyping. Computer, 22(5): 13–25, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:1989:OVC

[307] R. Gupta, W. H. Cheng, R. Gupta, I. Hardonag, and M. A. Breuer. An object-oriented VLSI CAD framework. Computer, 22(5):28–37, May 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jordan:1989:SSC

[308] P. W. Jordan, K. S. Keller, R. W.

Tucker, and D. Vogel. Software storming: Combining rapid prototyping and knowledge engineering. Computer, 22 (5):39–48, May 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1989:PSU

[309] T. G. Lewis, F. Handloser, III, S. Bose, and S. Yang. Prototypes from Standard User Interface Management Systems. Computer, 22(5):51–60, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Duke:1989:RPF

[310] E. L. Duke, R. W. Brumbaugh, and J. D.

Disbrow. A rapid prototyping facility for flight research in advanced systems concepts. Computer, 22(5):61–66, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cieslak:1989:PNP

[311] R. Cieslak, A. Fawaz, S. Sachs, P. Varaiya, J. Walrand, and A. Li. The programmable network prototyping system. Computer, 22(5):67–76, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McKeeman:1989:GTA

[312] W. M. McKeeman. Graduation talk at Wang Institute. Computer, 22(5): 78–80, May 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Iyengar:1989:AIM

[313] S. S. Iyengar and R. L. Kashyap. Autonomous intelligent machines. Computer, 22(6):14–17, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bares:1989:AAR

[314] John Bares, Martial Hebert, Takeo Kanade, Eric Krotkov, Tom Mitchell, Reid Simmons, and William Whittaker. Ambler — an autonomous rover for planetary exploration. Computer, 22(6): 18–26, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weisbin:1989:AMR

[315] C. R. Weisbin, G. de Saussure, J. R. Einstein, F. G. Pin, and E. Heer. Autonomous mobile robot navigation and learning. Computer, 22(6):29–35, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rao:1989:AFL

[316] N. S. V. Rao. Algorithmic framework for learned robot navigation in unknown terrains. Computer, 22(6):37–43, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Elfes:1989:UOG

[317] A. Elfes. Using occupancy grids for mobile robot perception and navigation.

Computer, 22(6):46–57, June 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bihari:1989:CAS

[318] T. E. Bihari, T. M. Walliser, and M. R.

Patterson. Controlling the adaptive suspension vehicle. Computer, 22(6): 59–65, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barhen:1989:NLC

[319] Jacob Barhen, Sandeep Gulati, and Michail Zak. Neural learning of constrained nonlinear transformations. Computer, 22(6):67–76, June 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1989:PRK

[320] I. Lee, R. B. King, and R. P. Paul. A predictable real-time kernel for distributed multisensor systems. Computer, 22(6): 78–83, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roth-Tabak:1989:BEM

[321] Y. Roth-Tabak and R. Jain. Building an environment model using depth information. Computer, 22(6):85–90, June 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trivedi:1989:VSR

[322] M. M. Trivedi, C. X. Chen, and S. B.

Marapane. A vision system for robotic inspection and manipulation. Computer, 22(6):91–97, June 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Preston:1989:ACB

[323] K. Preston, Jr. The Abingdon Cross benchmark survey. Computer, 22(7): 9–18, July 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1989:DIS

[324] J. E. Smith. Dynamic instruction scheduling and the Astronautics ZS-1. Computer, 22(7):21–35, July 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Biggerstaff:1989:DRM

[325] T. J. Biggerstaff. Design recovery for maintenance and reuse. Computer, 22 (7):36–49, July 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chisvin:1989:CAM

[326] L. Chisvin and R. J. Duckworth. Content-addressable and associative memory: Alternatives to the ubiquitous RAM. Computer, 22(7):51–64, July 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dinning:1989:SSM

[327] A. Dinning. A survey of synchronization methods for parallel computers. Computer, 22(7):66–77, July 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFanti:1989:VES

[328] T. A. DeFanti, M. D. Brown, and B. H. McCormick. Visualization: Expanding scientific and engineering research opportunities. Computer, 22(8):12–16, 22– 25, August 1989. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Helman:1989:RDV

[329] J. Helman and L. Hesselink. Representation and display of vector field topology in fluid flow data sets. Computer, 22(8): 27–36, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Long:1989:ARD

[330] M. B. Long, K. Lyons, and J. K. Lam.

Acquisition and representation of 2D and 3D data from turbulent flows and flames. Computer, 22(8):39–45, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuchs:1989:IVM

[331] Henry Fuchs, Marc Levoy, and Stephen M. Pizer. Interactive visualization of 3-D medical data. Computer, 22(8):46–51, August 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hibbard:1989:VLD

[332] W. Hibbard and D. Santek. Visualizing large data sets in the Earth sciences. Computer, 22(8):53–59, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kluksdahl:1989:RVS

[333] N. C. Kluksdahl, A. M. Kriman, and D. K. Ferry. The role of visualization in the simulation of quantum electronic transport in semiconductors. Computer, 22(8):60–66, August 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Phillips:1989:DVA

[334] R. L. Phillips. Distributed visualization at Los Alamos National Laboratory. Computer, 22(8):70–77, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cabral:1989:VTA

[335] B. Cabral and C. L. Hunter. Visualization tools at Lawrence Livermore National Laboratory. Computer, 22(8):77– 84, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haber:1989:SVR

[336] R. B. Haber. Scientific visualization and the Rivers Project at the National Center for Supercomputing Applications. Computer, 22(8):84–89, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bancroft:1989:SVC

[337] G. V. Bancroft, T. Plessel, F. Merritt, P. P. Walataka, and V. Watson. Scientific visualization in computational aerodynamics at NASA Ames Research Center. Computer, 22(8):89–95, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rosenblum:1989:VED

[338] L. J. Rosenblum. Visualization of experimental data at the Naval Research Laboratory. Computer, 22(8):95– 101, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1989:TIP

[339] T. L. Pappas. TEXnology on the IBM PC. Computer, 22(8):111–120, August 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1989:XSR

[340] J. Johnson, T. L. Roberts, W. Verplank, D. C. Smith, C. H. Irby, M. Beard, and K. Mackey. The Xerox Star: a retrospective. Computer, 22(9):11–26, 28– 29, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheng:1989:VPC

[341] H. Cheng. Vector pipelining, chaining, and speed on the IBM 3090 and Cray X-MP. Computer, 22(9):31–42, 44, 46, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lieberherr:1989:AGS

[342] Karl J. Lieberherr and Ian M. Holland. Assuring good style for objectoriented programs. Computer, 22(9):38– 44, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Artsy:1989:DPM

[343] Yeshayahu Artsy and Raphael Finkel.

Designing a process migration facility: The Charlotte experience. Computer, 22(9):47–56, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Review: Computing Reviews, Vol. 31, No. 8, August 1990.

Ford:1989:MSE

[344] G. A. Ford and N. E. Gibbs. A Master of Software Engineering Curriculum: Recommendations from the Software Engineering Institute. Computer, 22(9):59– 71, September 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ambler:1989:IVT

[345] A. L. Ambler and M. M. Burnett. Influence of visual technology on the evolution of language environments. Computer, 22(10):9–22, October 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roman:1989:DAV

[346] G.-C. Roman and K. C. Cox. A declarative approach to visualizing concurrent computations. Computer, 22(10):25– 36, October 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lehr:1989:VPD

[347] Ted Lehr, Zary Segall, Dalibor F. Vrsalovic, Eddie Caplan, Alan L. Chung, and Charles E. Fineman. Visualizing performance debugging. Computer, 22(10):38– 51, October 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kramer:1989:GCP

[348] Jeffrey Kramer, J. Magee, and K. Ng. Graphical configuration programming. Computer, 22(10):53–58, 62–65, October 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gu:1989:SAV

[349] J. Gu and K. F. Smith. A structured approach for VLSI circuit design. Computer, 22(11):9–22, November 1989. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hac:1989:SON

[350] A. Hac and H. B. Mutlu. Synchronous optical network and broadband ISDN protocols. Computer, 22(11):26–34, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singhal:1989:DDD

[351] M. Singhal. Deadlock detection in distributed systems. Computer, 22(11):37– 48, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gries:1989:TSR

[352] D. Gries and D. Marsh. The 1987–88 Taulbee Survey Report. Computer, 22 (11):49–56, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1989:CSP

[353] Anonymous. The computer spectrum: a perspective on the evolution of computing. Computer, 22(11):57–63, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1989:SHU

[354] Anonymous. Supercomputer hardware:

an update of the 1983 report’s summary and tables. Computer, 22(11):63– 68, November 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grosky:1989:GEI

[355] W. I. Grosky and R. Mehrotra. Guest Editors’ introduction: Image database management. Computer, 22(12):7–8, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kasturi:1989:MDP

[356] R. Kasturi, R. Fernandez, M. L. Amlani, and W.-C. Feng. Map data processing in Geographic Information Systems. Computer, 22(12):10–21, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brolio:1989:IDS

[357] J. Brolio, B. A. Draper, J. R. Beveridge, and A. R. Hanson. ISR: A database for symbolic processing in computer vision. Computer, 22(12):22–30, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jagadish:1989:OMI

[358] H. V. Jagadish and L. O’Gorman. An object model for image recognition. Computer, 22(12):33–41, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goodman:1989:KBC

[359] Andrew M. Goodman, Robert M. Haralick, and Linda G. Shapiro. Knowledgebased computer vision — integrated programming language and data management system design. Computer, 22(12): 43–54, December 1989. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pizano:1989:SSI

[360] A. Pizano, A. Klinger, and A. Cardenas. Specification of spatial integrity constraints in pictorial databases. Computer, 22(12):59–71, December 1989. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ziv:1977:UAS

[361] Jacob Ziv and Abraham Lempel. A universal algorithm for sequential data compression. IEEE Transactions on Information Theory, IT-23(3):337–343, May 1977. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Ziv:1978:CIS

[362] J. Ziv and A. Lempel. Compression of individual sequences via variable rate coding. IEEE Transactions on Information Theory, IT-24(5):530–536, September 1978. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Crocker:1995:PPN

[363] Lee Daniel Crocker. PNG: The Portable Network Graphic format. Dr. Dobb’s Journal of Software Tools, 20(7):36, 38, 40, 42, 44, July 1995. CODEN DDJOEB. ISSN 1044-789X.

Roelofs:1999:PDG

[364] Greg Roelofs. PNG: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-542-4. xix + 327 + 4 pp. LCCN T385 .R58 1999. US$32.95. URL http://www.

libpng.org/pub/png/pngbook.html;
 http://www.oreilly.com/catalog/
pngdefg.

Salomon:2000:DCC

[365] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-95045-1. xvi + 823 pp. LCCN QA76.9.D33 S25 2000. US$39.95. URL http://

                                        www.ecs.csun.edu/~dxs/DCadvertis/
                                        Dcomp2Ad.html.